OpenStudy (anonymous):

Rewrite with only sin x and cos x. sin 3x - cos x

OpenStudy (callisto):

sin 3x - cos x = sin(2x+x) -cos x Use sin(a+b) = sina cosb + sinb cosa = sin(2x)cos(x) + cos(2x)sin(x) - cosx sin(2x) = 2sin(x)cos(x) cos(2x) = 1-2(sinx)^2 Sub. these two identities into sin(2x)cos(x) + cos(2x)sin(x) - cosx Let see what you get..

OpenStudy (anonymous):

2 sin x cos2x + sin x - 2 sin3x - cos x right?

OpenStudy (callisto):

Yes.

OpenStudy (anonymous):

Really? Thanks!

OpenStudy (callisto):

Really :)